# Exploring Adapter Design Tradeoffs for Low Resource Music Generation

### v1

@@ 4.2. Subjective Evaluations
Across both genres, linear adapters (MG-L) consistently underperformed in subjective evaluations, aligning with objective metrics where they recorded higher FAD and FD scores, especially at larger parameter scales. A key insight from our analysis is the contrast between objective and subjective performance, particularly in the case of Hindustani Classical music. Here, Mustango models achieved lower FAD scores, indicating greater diversity and stronger adherence to prompt content in terms of entropy. However, this did not translate to better human evaluations—annotators found that despite its creative variety, Mustango often lacked note alignment, structural stability, and audio clarity, which ultimately impacted the perceived quality of the generated music. In contrast, MusicGen-CNN (MG-C), though slightly less diverse, was rated higher for coherence, clarity, and musicality explaining the difference in objective and subjective evaluations.
@@ 4.3. Computational Efficiency
  • MusicGen shows excellent efficiency in its top-performing adapters: for Hindustani Classical, the 40M Convolution (MG-C) configuration takes only 20 hours of GPU time, while for Turkish Makam, the 40M Transformer (MG-T) configuration takes 22 hours.
  • Mustango, which benefits most from a Convolution-based adapter for both Hindustani and Makam at the 40M scale (MS-C), requires about 45 hours of GPU time. Although longer than its MusicGen counterpart, it is still substantially more efficient than training a large model from scratch.
@@ 5. Conclusion and Future Work
Finally, our results suggest that the 40M parameter scale is well-suited for the dataset sizes used in this study, although optimal adapter sizes may vary with larger or more complex datasets. While our study provides valuable insights into adapter-based PEFT for music generation, several directions remain open for future exploration:
  • Hybrid PEFT Approaches: Investigating hybrid approaches that combine adapters with LoRA (Hu et al., 2022) or prefix tuning (Li and Liang, 2021) could provide insights into maximizing both efficiency and adaptability, potentially leading to even more lightweight and effective fine-tuning techniques for music generation models.
  • Expanding to Other Musical Traditions: This study focused on Hindustani Classical and Turkish Makam music, but similar methods could be applied to Carnatic (Vijayakrishnan, 2007), Persian Dastgah (Nettl, 2001), Gamelan (Becker, 1993), and other culturally significant music traditions to assess how adapter-based fine-tuning generalizes across diverse genres.
  • Exploring Cross-Model Adapter Transfer: Investigating whether adapters fine-tuned on transformer-based models (MusicGen) can be transferred and adapted for diffusion-based models (Mustango). This could enable cross-architecture PEFT (Yang et al., 2024), making fine-tuning more generalizable across different generative paradigms.
  • Identifying Data vs Model Size Tradeoffs: Similar to NLP, music generation models show sensitivity to dataset and model size trade-offs. Inspired by Chinchilla (Hoffmann et al., 2022) and other scaling studies (Kaplan et al., 2020), future work could explore scaling laws for music generation, analyzing how performance varies with dataset size, model capacity, and genre diversity.

### v2

@@ 4.2. Subjective Evaluations
Across both genres, linear adapters (MG-L) consistently underperformed in subjective evaluations, aligning with objective metrics where they recorded higher FAD and FD scores, especially at larger parameter scales. A key insight from our analysis is the contrast between objective and subjective performance, particularly in the case of Hindustani Classical music. Here, Mustango models achieved lower FAD scores, indicating stronger adherence to prompt which implies that it is closer to the ground truth distribution in terms of entropy. However, this did not translate to better human evaluations—annotators found that despite its creative variety, Mustango often lacked note alignment, structural stability, and audio clarity, which ultimately impacted the perceived quality of the generated music. In case of diversity, we observed that the MusicGen model generations were highly repetitive in its generation and had very narrow set of instruments, rhythm types and melodies in its generations leading to our understanding that Mustango could generate music using a broader set of musical attributes including instruments, rhythms and melodies which led to a lower FAD score since it is closer to the ground truth distribution. In contrast, MusicGen-CNN (MG-C), though slightly less diverse, was rated higher for coherence, clarity, and musicality explaining the difference in objective and subjective evaluations.
@@ 4.3. Computational Efficiency
  • MusicGen shows excellent efficiency in its top-performing adapters: for Hindustani Classical, the 40M Convolution (MG-C) configuration takes only 20 hours of GPU time, while for Turkish Makam, the 40M Transformer (MG-T) configuration takes 22 hours. For MusicGen configurations, at inference time, generating each sample requires approximately 3 seconds with 40GB of GPU memory.
  • Mustango, which benefits most from a Convolution-based adapter for both Hindustani and Makam at the 40M scale (MS-C), requires about 45 hours of GPU time. Although longer than its MusicGen counterpart, it is still substantially more efficient than training a large model from scratch. At inference time, Mustango requires 100 seconds (depending on the number of denoising steps which in our case was 200) for generating 10 second audios with a batch size of 4 and GPU memory of 32GB on a single GPU.
@@ 5. Conclusion and Future Work
Our results indicate that the 40M parameter scale is well-suited for the dataset sizes considered, though optimal adapter configurations may vary with larger or more complex datasets. While our study offers key insights into adapter-based PEFT for music generation, several promising directions remain:
  • Hybrid PEFT Methods: Combining adapters with techniques like LoRA (Hu et al., 2022) or prefix tuning (Li and Liang, 2021) may enhance efficiency and adaptability.
  • Cultural Extension: Applying our approach to traditions like Carnatic (Vijayakrishnan, 2007), Persian Dastgah (Nettl, 2001), or Gamelan (Becker, 1993) can test generalizability across diverse genres.
  • Cross-Architecture Transfer: Exploring whether adapters trained on transformer-based models (e.g., MusicGen) can be transferred to diffusion-based ones (e.g., Mustango) (Yang et al., 2024).
  • Scaling Trade-offs: Inspired by Chinchilla (Hoffmann et al., 2022) and scaling laws (Kaplan et al., 2020), future work can investigate the interplay between model size, data volume, and genre diversity.